Solve the triangle, a=6,b=8,C=70°
Answer by Boreal(15235)

Law of cosines
a=6
b=8
C=70 degrees
c^2=a^2+b^2-2ab cos C
c^2=48+64-2(48)(.3420)
c^2=112-32.83=79.17
c=8.90